# Impact of heterogeneity on infection probability: Insights from single-hit dose-response models

### v1

@@ 4.5.2 Example $C_{2}^{\prime}$: Beta-distributed infectivity and negative binomial dose distribution
For a fixed mean dose, the expected probability of invasion of model $C_{2}^{\prime}$ is a decreasing function of the dose variance.
@@
The second inequality stems from the non-convexity of $P^{(C_{1})}(\mu_{x},v_{x},n)$ with respect to $n$. Applying Jensen’s inequality, $P^{(C_{1})}(\mu_{x},v_{x},{\text{E}}_{n}(n))\geq{\text{E}}_{n}(P^{(C_{1})}(\mu_{x},v_{x},n))$ is derived. This reduces to $P^{(C_{1})}(\mu_{x},v_{x},\mu_{n})\geq P^{(C_{1}^{\prime})}(\mu_{x},v_{x},\mu_{n})$ using $\mu_{n}={\text{E}}_{n}(n)$ and $P^{(C_{1}^{\prime})}(\mu_{x},v_{x},\mu_{n})={\text{E}}_{n}(P^{(C_{1})}(\mu_{x},v_{x},n))$ (cf. Eq. (44)).
@@
Figure 6: Graphical illustration of the inequalities given in Proposition 4.5. The dependence of the probability of infection on the mean infectivity for a system with homogeneous infectivity (Model $A$, continuous black line) is compared to the dependence on $\mu_{x}$ of the expected probability of infection of models $C_{1}$ (dashed red), $C^{\prime}_{1}$ (dot-dashed blue) and $C^{\prime}_{2}$ (long dashed green). The dose was set to $n=5$ for model $A$. The mean and variance of the dose were set to $\mu_{n}=5$ and $v_{n}=20$, respectively, for models $C_{1}$, $C^{\prime}_{1}$ and $C^{\prime}_{2}$.
5 A population growth model with random between-host infectivity
This section provides a biologically explicit interpretation of single-hit dose-response models of type $C$ in which the between-host heterogeneity of infectivity is described as a deterministic population growth process with heterogeneity between hosts.
@@ 5.0.1 General formulation of the model
This section derives a biologically motivated probability distribution for the infectivity based on a simple growth model [37] for the microbial population within a host. In particular, it is assumed that microbes duplicate at a rate $\beta$ and die at a rate $\delta$. At the early stages of microbial growth, the number $N$ of microbes is assumed to obey the equation $\text{d}N/\text{d}t=(\beta-\delta)N$. A necessary condition for a microbial population to grow within a host is that the basic reproductive ratio, $R=\beta/\delta$, is larger than 1. In this case, microbes produce more than one offspring on average during their average lifetime, $1/\delta$. In contrast, a microbial population dies out when $R<1$.
Let us assume that the basic reproductive ratio $R_{i,h}$ of a microbe $i$ in host $h$ is a random variable with PDF $\rho_{R}(R_{i,h};\bm{\xi}_{h})$, where $\bm{\xi}_{h}$ is a set of host-dependent parameters. The infectivity of a randomly chosen microbe within host $h$ is the probability that such microbe grows within the host, i.e. the probability that $R_{i,h}>1$:
@@ 5.0.2 Example: Linking heterogeneous infectivity and microbial birth and death
As expected, the mean infectivity of the microbe, $\mu_{x}$, decreases with $\mu_{\lambda}$ (see the contour levels in Fig. 7(a)), i.e. higher mortality of the microbe leads to lower infectivity. Conversely, there is a positive association between $\mu_{x}$ and $v_{\lambda}$, indicating that greater variability in microbe mortality tends to enhance infectivity on average.
Regions characterised by low infectivity variance, $v_{x}$, are evident in regimes (i) and (vi), visible as dark areas in Fig. 7(b). In these contexts, models featuring homogeneous infectivity (such as type $A$ and $A^{\prime}$) can offer a reasonable description of the expected probability of infection. In regime (i), the low infectivity variance region corresponds to situations with high mean infectivity, which forces the maximum infectivity variance $v_{\max}(\mu_{x})$ to be small. Conversily, regime (iv) corresponds to systems characterised by relatively homogeneous infectivity, irrespectively of the mean infectivity.
@@ 6 Conclusion
This work has delved into the impact of variation in dose and microbial infectivity on infection probability within the framework of single-hit dose-response models. It has been rigorously shown that heterogeneity in the infectivity of microbial populations ingested by a host enhances the chances of infection (heterogeneity of type I, Theorem 2.2). This finding should be differentiated from the observation that the expected probability of infection solely depends on the mean infectivity, $\mu_{m}$. In essence, the expected probability represents the average infection likelihood across a group of hosts exposed to doses with heterogeneous infectivity, rather than the infection probability of an individual.
The presence of heterogeneity of type II, where microbial infectivity varies across hosts, has been demonstrated to reduce the expected probability of infection. General results have been given for microbial populations with small infectivity (Theorem 4.1). Moreover, the ubiquitous flattening of the dose-response curve observed in experiments has also been shown to be associated with an increase in the variance of infectivity. The validity of these results has been illustrated with several models which assume random infectivity between hosts. Across these models, trends indicating heightened infectivity variance remain consistent regardless of the infectivity expectation, as summarised in Proposition 4.5. This suggests that the small infectivity limit required in Theorem 4.1 may not be overly restrictive. Investigating a more general theorem which relaxes this condition might be interesting.
@@
The last results section has proposed a microbial growth model with randomly distributed reproductive ratio within a host. An example of this model has been presented, utilising the effective microbial death rate as a parameter, which essentially acts as the inverse of the reproductive ratio. The model provides interesting predictions: near-uniform infectivity is anticipated when either the mean or variance of the microbial death rate is small. Conversely, systems characterised by substantial mean and variance in the effective death rate will exhibit heterogeneous infectivity. Furthermore, this model offers a biological interpretation for the parameters of the beta-Poisson dose-response model (referred to as Example $C_{1}^{\prime}$ within the framework presented here). Specifically, it has been found that both $\alpha$ and $\beta$ decrease with increasing heterogeneity in the effective microbial death rate. In contrast, both $\alpha$ and $\beta$ increase when increasing the expectation of the effective microbial death rate.
@@ Acknowledgements
The author acknowledges fruitful discussions with Norval Strachan and Ovidiu Rotariu as well as funding support from a Medical Research Council Fellowship (MR/W021455/1).

### v2

@@ 4.5.2 Example $C_{2}^{\prime}$: Beta-distributed infectivity and negative binomial dose distribution
For a fixed mean dose, the expected probability of infection of model $C_{2}^{\prime}$ is a decreasing function of the dose variance.
@@
The second inequality stems from the non-convexity of $P^{(C_{1})}(\mu_{x},v_{x},n)$ with respect to $n$. Applying Jensen’s inequality, leads to $P^{(C_{1})}(\mu_{x},v_{x},{\text{E}}_{n}(n))\geq{\text{E}}_{n}(P^{(C_{1})}(\mu_{x},v_{x},n))$. This reduces to $P^{(C_{1})}(\mu_{x},v_{x},\mu_{n})\geq P^{(C_{1}^{\prime})}(\mu_{x},v_{x},\mu_{n})$ using $\mu_{n}={\text{E}}_{n}(n)$ and $P^{(C_{1}^{\prime})}(\mu_{x},v_{x},\mu_{n})={\text{E}}_{n}(P^{(C_{1})}(\mu_{x},v_{x},n))$ (cf. Eq. (46)).
@@
Figure 6: Graphical illustration of the inequalities given in Proposition 4.5. The dependence of the probability of infection on the mean infectivity for a system with homogeneous infectivity (Model $A$, continuous line) is compared to the dependence on $\mu_{x}$ of the expected probability of infection of models $C_{1}$ (dashed line), $C^{\prime}_{1}$ (dot-dashed line) and $C^{\prime}_{2}$ (long dashed line). The dose was set to $n=5$ for model $A$. The mean and variance of the dose were set to $\mu_{n}=5$ and $v_{n}=20$, respectively, for models $C_{1}$, $C^{\prime}_{1}$ and $C^{\prime}_{2}$.
5 A model for microbial infection with random infectivity
This section provides a biologically explicit interpretation of single-hit dose-response models of type $C$, which account for variations in pathogen infectivity across different hosts. Between-host heterogeneity in infectivity is assumed to arise from biological factors that influence the ability of microbes to establish an infection. In line with Assumption 1 in the introduction, the proposed framework allows for infection to be initiated by a single microbe but specifically considers cases where microbial growth is essential for infection establishment.
Inspired by compartmental models of infection dynamics within a host, it is assumed that infection occurs if the microbial reproduction number within the host, $R$, satisfies $R>1$ [37, 38, 39]. Conversely, if $R<1$, the infection fails to establish. For instance, in a minimal within-host model, microbial dynamics can be described by a simple growth mechanism where microbes infect cells at rate $\beta$ and die at rate $\delta$  [40, 41]. At the early stages of infection, the microbial population $N$ is assumed to obey the equation $\text{d}N/\text{d}t=(\beta-\delta)N$. A necessary condition for infection to occur is $R=\beta/\delta>1$, meaning each microbe produces more than one offspring on average during its average lifetime, $1/\delta$. In contrast, a microbial population dies out when $R<1$. While more complex within-host models incorporate additional parameters, the fundamental infection criterion $R>1$ remains broadly applicable [37, 38, 39].
@@ 5.0.1 General formulation of the model
Between-host heterogeneity is modelled by assuming that the basic reproductive ratio $R_{i,h}$ of a microbe $i$ in host $h$ is a random variable with PDF $\rho_{R}(R_{i,h};\bm{\xi}_{h})$, where $\bm{\xi}_{h}$ is a set of host-dependent parameters. The infectivity of a randomly chosen microbe within host $h$ is the probability that $R_{i,h}>1$:
@@ 5.0.2 Example: Linking heterogeneous infectivity and microbial birth and death
As expected, the mean infectivity of the microbe, $\mu_{x}$, decreases with $\mu_{\lambda}$ (see the contour levels in Fig. 7(a)), i.e. higher mortality of the microbe leads to lower infectivity. Conversely, there is a positive association between $\mu_{x}$ and $v_{\lambda}$, indicating that greater variability in microbe mortality tends to enhance infectivity on average. This enhanced infectivity in populations with heterogeneous mortality rates is likely driven by microbes with lower-than-average $\lambda$, which survive longer and are more likely to infect the host than microbes in a homogeneous population with the same mean mortality $\mu_{\lambda}$.
Regions characterised by low infectivity variance, $v_{x}$, are evident in regimes (i) and (vi), visible as dark areas in Fig. 7(b). In these contexts, models featuring homogeneous infectivity (such as type $A$ and $A^{\prime}$) can offer a reasonable description of the expected probability of infection. In regime (i), the low infectivity variance region corresponds to situations with high mean infectivity, which forces the maximum infectivity variance $v_{\max}(\mu_{x})$ to be small (see Eq. (34)). Conversily, regime (iv) corresponds to systems characterised by relatively homogeneous infectivity, irrespectively of the mean infectivity.
@@ 6 Conclusion
This work has delved into the impact of variation in dose and microbial infectivity on infection probability within the framework of single-hit dose-response models. It has been rigorously shown that heterogeneity in the infectivity of microbial populations ingested by a host enhances the chances of infection (heterogeneity of type I, Theorem 2.2). This finding should be differentiated from the observation that the expected probability of infection solely depends on the mean infectivity, $\mu_{m}$ (Sec. 4.3). In essence, the expected probability represents the average infection likelihood across a group of hosts exposed to doses with heterogeneous infectivity, rather than the infection probability of an individual.
The presence of heterogeneity of type II, where microbial infectivity varies across hosts, has been demonstrated to reduce the expected probability of infection. General results have been given for microbial populations with small infectivity (Theorem 4.1). Moreover, the ubiquitous flattening of the dose-response curve observed in experiments has also been shown to be associated with an increase in the variance of infectivity. The validity of these results has been illustrated with several models which assume random infectivity between hosts. Across these models, the expected probability of infection decreases when increasing the heterogeneity of infectivity regardless of the infectivity expectation, as summarised in Proposition 4.5 (cf. $P^{(A)}$ and $P^{(C_{1})}$). This suggests that the small infectivity limit required in Theorem 4.1 may not be overly restrictive. Investigating a more general theorem which relaxes this condition might be interesting.
@@
The last results section has proposed a within-host microbial growth model with randomly distributed reproductive ratio within a host. An example of this model has been presented, utilising the effective microbial death rate as a parameter, which essentially acts as the inverse of the reproductive ratio. The model provides interesting predictions: near-uniform infectivity is anticipated when either the mean or variance of the microbial death rate is small. Conversely, systems characterised by substantial mean and variance in the effective death rate will exhibit heterogeneous infectivity. Furthermore, this model offers a biological interpretation for the parameters of the beta-Poisson dose-response model (referred to as Example $C_{1}^{\prime}$ within the framework presented here). Specifically, it has been found that both $\alpha$ and $\beta$ decrease with increasing heterogeneity in the effective microbial death rate. In contrast, both $\alpha$ and $\beta$ increase when increasing the expectation of the effective microbial death rate.
@@ Acknowledgements
The author acknowledges fruitful discussions with Norval Strachan, Ovidiu Rotariu and Ken Forbes as well as funding support from a Medical Research Council Fellowship (MR/W021455/1).